# DextrAH-G: Pixels-to-Action Dexterous Arm-Hand Grasping with Geometric Fabrics

### v2

@@ Abstract
Figure 1: DextrAH-G continuously controls a dexterous robot to grasp and transport a diverse range of objects directly from streaming depth images.
@@ 1 Introduction
Grasping, a fundamental skill coarsely mastered by even two-year-olds [1], still poses a challenge in robotics. Equipping machines with proficient grasping skill is paramount to automating aspects of logistics, manufacturing, space, and search-and-rescue among many others use cases. To date, robot grasping has emerged as a chief research problem with many advancements made. However, existing methods often have limited speed, dexterity, and reliability, especially with high-actuator-count embodiments. Moreover, partial observability and succinct coordination over a large number of motors in dexterous platforms pose additional challenges in acquiring proficient grasping skill.
@@
To address these challenges with real-world dexterous robot grasping, we propose DextrAH-G: a combined pixels-to-action fabric-guided policy (FGP) and geometric fabric controller that achieves state-of-the-art grasping performance in the real world. The main contributions of this work include: 1) a vectorized geometric fabric controller that creates an inductive bias for policy learning, avoids collision, upholds joint constraints, and shapes the behavior, 2) simulation-only RL training of a privileged FGP over vectorized geometric fabrics that enables high-performance, multi-object grasping, 3) depth-based, multi-modal FGP distillation of the privileged FGP that replicates the original behavior and enables object position predictions and 4) zero-shot sim2real transfer of DextrAH-G with new state-of-the-art dexterous grasping performance on many diverse novel objects in the real world, a significant progression towards grasp-anything skill in dexterous robotics.
@@ 2.2 Policy Learning and Control
Learned policies for robots typically issue actions to simple controllers like joint-level proportional-derivative (PD) control or operational space control (OSC). Consequently, controller simplicity shifts the burden of discovering and imparting well-controlled behavior entirely to the learned policy itself. Globally, well-controlled behavior has complex, multi-faceted priorities such as primary objective completion, various hardware constraints, collision avoidance, and other qualities like natural movements. Discovering this rich behavioral conglomerate via learning is very challenging due to optimization locality, approximation and generalization errors inherit to neural networks, behavior specification, exploration, and high-dimensional action spaces. This arduous path can be avoided by embedding many of these behaviors within more sophisticated control layers. Recently, fabric-guided policies (FGP) were trained over a geometric fabric via RL in simulation to achieve new state-of-the-art in-hand cube reorientation performance in the real world [13]. The geometric fabric itself handled joint hardware constraints, created an input action space for controlling the hand fingertips, and guided the fingertips towards making contact with the cube. Leveraging such a rich controller promoted a much simpler reward function predominantly geared towards primary task completion. Other works also mix sophisticated control and policy learning like Riemannian Motion Policies in [14] and QP control in [15]. Since geometric fabrics have been shown to outperform RMPs [16, 17], DMPs [17], Koopman Operator policies [18], and derive from strong theoretical analyses [19, 16, 20], we elect to use geometric fabrics within this work to construct DextrAH-G.
3 DextrAH-G
@@
Figure 2: Our approach trains a teacher privileged fabrics-guided policy (FGP) using RL (top), distills the privileged FGP into a student depth FGP to predict the teacher’s actions and object position (middle), and deploys DextrAH-G with a state machine in the real world for bin packing (bottom).
@@ 3.1 Geometric Fabrics and Fabric-Guided Policies (FGPs)
where $\mathbf{M}_{f}\in\mathbb{R}^{n\times n}$ is the positive-definite system metric (mass), which captures system prioritization, $\mathbf{f}_{f}\in\mathbb{R}^{n}$ is a nominal path generating geometric force, and $\mathbf{f}_{\pi}(\mathbf{a})\in\mathbb{R}^{n}$ is an additional driving force of some action $\mathbf{a}\in\mathbb{R}^{m}$. $\mathbf{q}_{f},{\dot{\mathbf{q}}}_{f},{\ddot{\mathbf{q}}}_{f}\in\mathbb{R}^{n}$ are the position, velocity, and acceleration of the fabric. This fundamental equation produces an acceleration ${\ddot{\mathbf{q}}}_{f}$, which evolves the fabric state $\mathbf{q}_{f}$ and ${\dot{\mathbf{q}}}_{f}$ over time through numerical integration. One can see that $\mathbf{f}_{\pi}$ influences ${\ddot{\mathbf{q}}}_{f}$, and thereby, the fabric state. For in-depth discussion of geometric fabrics, we refer the reader prior works [19, 16, 20, 13]. See Appendix B.1 for details on how a geometric fabric connects with a robot.
@@
We create a geometric fabric as part of DextrAH-G for three primary reasons: 1) upholding hardware safety constraints, 2) creating an inductive bias via the exposed action space that simultaneously guides policy exploration and favorably shapes the overall robot motion, and 3) maintaining the robot posture to promote kinematic manipulability (see Appendix B.3 for details on posture control).
Hardware Safety: The two primary hardware safety concerns during real-world deployment are respecting joint constraints and avoiding undesirable collisions. We address joint constraints via fabric design as in [13] (see Appendix B.2). Our design for collision avoidance is as follows.
@@
Both environmental and self-collision avoidance are explicitly handled by the fabric. This behavior is constructed by first approximating the geometry of the robot as a collection of spheres. We use the forward kinematics of the arm to map robot configuration to the origin of every sphere attached to the robot body, $\mathbf{x}=\phi_{fk}(\mathbf{q})$. We then construct a base metric response per point as $\mathbf{M}_{b}=\sum_{i}\frac{s_{i}}{\underline{d}_{i}}\hat{\mathbf{n}_{i}}\otimes\hat{\mathbf{n}_{i}}$, where $s_{i}=\frac{1}{2}\tanh(-\alpha_{1}(v_{i}-\alpha_{2})+1)$ is a smooth velocity gate that goes high when this sphere point is moving towards collision body $i$ (can be any known collision object in the scene and other body spheres). $v_{i}=-{\dot{\mathbf{x}}}\cdot\hat{\mathbf{n}}_{i}$ is the signed impact speed that is positive when moving away from collision and negative towards collision. $\hat{\mathbf{n}}_{i}=\frac{\mathbf{r}_{i}-\mathbf{x}}{\|\mathbf{r}_{i}-\mathbf{x}\|}$ is the direction from the sphere point to the closest point on collision body $i$, $\mathbf{r}_{i}$. Both $\alpha_{1},\alpha_{2}\in\mathbb{R}^{+}$ are gains. Additionally, $\underline{d}_{i}=\max(d_{min},d_{i})$ is a positively lower-bounded distance, where $d_{min}\in\mathbb{R}^{+}$ and $d_{i}$ is the signed distance between the body sphere and collision body $i$. Finally, we normalize this metric to maintain its Eigenspectrum (directions of importance) while maintaining constant norm invariant to the number of collision bodies, i.e., $\hat{\mathbf{M}}_{b}=\frac{\mathbf{M}_{b}}{\|\mathbf{M}_{b}\|}$. Next, we construct a base acceleration response per sphere point as ${\ddot{\mathbf{x}}}_{b}=-\sum_{i}\frac{1}{\underline{d}_{i}}\hat{\mathbf{n}_{i}}$, causing acceleration away from collision along the collision direction. We also normalize this base acceleration, $\hat{{\ddot{\mathbf{x}}}}_{b}=\frac{{\ddot{\mathbf{x}}}_{b}}{\|{\ddot{\mathbf{x}}}_{b}\|}$. We use these base responses to construct a geometric fabric term and forcing fabric term. Both terms have metric design $\mathbf{M}=\frac{\beta}{\widetilde{d}^{2}}\hat{\mathbf{M}}_{b}$, where $\beta\in\mathbb{R}^{+}$ is a gain, and $\widetilde{d}=\min_{i}\{\underline{d}_{i}\}$. The geometric acceleration is ${\ddot{\mathbf{x}}}=k_{g}\|{\dot{\mathbf{x}}}\|^{2}\hat{{\ddot{\mathbf{x}}}}_{b}$, where $k_{g}\in\mathbb{R}^{+}$ is a gain. The forcing acceleration is ${\ddot{\mathbf{x}}}=k_{f}\hat{{\ddot{\mathbf{x}}}}_{b}-b{\dot{\mathbf{x}}}$, where $k_{f}\in\mathbb{R}^{+}$ is a gain and $b\in\mathbb{R}^{+}$ is a positive damping scalar. Overall, the geometric term is tuned to dominate the collision avoidance behavior with speed-invariant paths, while the forcing term prevents penetration near the collision boundary.
@@
Action Space: During grasping, human finger motions exhibit highly correlated movements that do not fully exercise the much higher degree of actuation afforded by our hands. Moreover, established grasps can most broadly be categorized into power and precision grasps [21]. Absent intricate in-hand manipulation, grasping motion can be concisely captured by movement on a lower dimensional manifold. We create such a manifold for the Allegro hand by retargeting human grasping motion data to the Allegro (see Appendix D for retargeting details) and apply principal component analysis to the motion dataset. Let $\mathbf{A}\in\mathbb{R}^{5\times 16}$ be the first five principal components from PCA and define $\widetilde{\mathbf{A}}=[\mathbf{0},\mathbf{A}]\in\mathbb{R}^{5\times 23}$, then the taskmap from the full robot configuration space to PCA space is $\mathbf{x}=\widetilde{\mathbf{A}}\mathbf{q}\in\mathbb{R}^{5}$. Within this taskmap, we define an attraction fabric term. The metric $\mathbf{M}(\mathbf{x})=m\mathbf{I}$ is a constant isotropic mass, where $m\in\mathbb{R}^{+}$. ${\ddot{\mathbf{x}}}=-k_{a}\tanh(\alpha_{a}\|\mathbf{x}-\mathbf{x}_{pca,target}\|)\frac{\mathbf{x}-\mathbf{x}_{pca,target}}{\|\mathbf{x}-\mathbf{x}_{pca,target}\|}-b{\dot{\mathbf{x}}}$, where $k_{a},\alpha_{a}\in\mathbb{R}^{+}$ are gains and $\mathbf{x}_{pca,target}$ is a target position in this space. Since this space is five dimensional, $\mathbf{x}_{pca,target}$ becomes the five dimensional action space for the hand. The attractor accelerates towards $\mathbf{x}_{pca,target}$ resulting in convergence towards these goals.
@@
To coordinate finger control with arm control, we create another action space that controls the pose of the palm. We accomplish this by creating a new taskmap that uses forward kinematics to map to 7 three-dimensional points attached to the palm, stacked into a 21 dimensional space. The attraction fabric in this space is the same as designed for the hand, but scales to the full 21 dimensions with the goal action position $\mathbf{x}_{g}\in\mathbb{R}^{21}$. To avoid a 21 dimensional action space for the arm, the overlying policy’s action space is only 6 dimensional and consists of three translational positions and three Euler angles for the palm. These quantities are transformed to 3D point targets for all 7 palm-fixed points, i.e., $\mathbf{x}_{g}$, and issued to the fabric. Across the full robot, the action space is 11 dimensional.
@@ 3.2 Teacher Privileged FGP Training (Reinforcement Learning)
We cast dexterous grasping as a reinforcement learning problem and train a privileged-state teacher policy in simulation to adeptly grasp 140 different objects (see Appendix E.8). Since the geometric fabric action space ensures that the robot will execute safe and natural behaviors, our reward design centers entirely around fingertip-object contact and lifting the object to a goal. This eases reward engineering and reward hacking behavior from the policy. We now cover training details below with additional details in Appendix E.
@@
Environment Modifications to Improve Robustness of Grasping Behavior: One of the key goals of this work is achieving robust grasping behavior. RL policies often converge to unnatural solutions that work well in simulation, but often fail to transfer to the real world (see Appendix E.13 for frail grasping behavior.) To learn robust grasping behavior, we introduce environment modifications as follows. Random Wrench Perturbations: We apply random wrenches that move and rotate the object in unpredictable ways (see Appendix E.3). This forces the policy to learn grasps that are robust to exogenous perturbations. Pose Noise: We add uncorrelated and correlated noise to the object pose observation (see Appendix E.4). This gives incentive to learn grasping behavior that opens the hand wider than typically needed when approaching the object to reduce unexpected contact and account for uncertainty in position and geometry. Friction Reduction: We reduce the default coefficient of friction of the object to $\mu=0.7$, which makes the objects more slippery, thereby, mitigating grasping behavior that is overly reliant on friction. Domain Randomization: We employ domain randomization across simulation parameters to learn policies that are robust across a spectrum of dynamics (see Appendix E.5).

### v3

@@ Abstract
Figure 1: DextrAH-G (Dexterous Arm-Hand Grasping) continuously controls a dexterous robot to grasp and transport a diverse range of objects directly from streaming depth images.
@@ 1 Introduction
Grasping, a fundamental skill coarsely mastered by even two-year-olds [1], still poses a challenge in robotics. Equipping machines with proficient grasping skill is paramount to automating aspects of logistics, manufacturing, space, and search-and-rescue among many other use cases. To date, robot grasping has emerged as a chief research problem with many advancements made. However, existing methods often have limited speed, dexterity, and reliability, especially with high-actuator-count embodiments. Moreover, partial observability and succinct coordination over a large number of motors in dexterous platforms pose additional challenges in acquiring proficient grasping skill.
@@
To address these challenges with real-world dexterous robot grasping, we propose DextrAH-G: a combined pixels-to-action fabric-guided policy (FGP) and geometric fabric controller that achieves state-of-the-art grasping performance in the real world. The main contributions of this work include: 1) a vectorized geometric fabric controller that creates an inductive bias for policy learning, avoids collision, upholds joint constraints, and shapes the behavior, 2) simulation-only RL training of a privileged FGP over vectorized geometric fabrics that enables high-performance grasping of many different objects, 3) depth-based, multi-modal FGP distillation of the privileged FGP that replicates the original behavior and enables object position predictions and 4) zero-shot sim2real transfer of DextrAH-G with new state-of-the-art dexterous grasping performance on many diverse novel objects in the real world, a significant progression towards grasp-anything skill in dexterous robotics.
@@ 2.2 Policy Learning and Control
Learned policies for robots typically issue actions to simple controllers like joint-level proportional-derivative (PD) control or operational space control (OSC). Consequently, controller simplicity shifts the burden of discovering and imparting well-controlled behavior entirely to the learned policy itself. Globally, well-controlled behavior has complex, multi-faceted priorities such as primary objective completion, various hardware constraints, collision avoidance, and other qualities like natural movements. Discovering this rich behavioral conglomerate via learning is very challenging due to optimization locality, approximation and generalization errors inherent to neural networks, behavior specification, exploration, and high-dimensional action spaces. This arduous path can be avoided by embedding many of these behaviors within more sophisticated control layers. Recently, fabric-guided policies (FGP) were trained over a geometric fabric via RL in simulation to achieve new state-of-the-art in-hand cube reorientation performance in the real world [13]. The geometric fabric itself handled joint hardware constraints, created an input action space for controlling the hand fingertips, and guided the fingertips towards making contact with the cube. Leveraging such a rich controller promoted a much simpler reward function predominantly geared towards primary task completion. Other works also mix sophisticated control and policy learning like Riemannian Motion Policies in [14] and QP control in [15]. Since geometric fabrics have been shown to outperform RMPs [16, 17], DMPs [17], Koopman Operator policies [18], and derive from strong theoretical analyses [19, 16, 20], we elect to use geometric fabrics within this work to construct DextrAH-G.
3 DextrAH-G: Dexterous Arm-Hand Grasping
@@
Figure 2: We train a privileged fabrics-guided policy (FGP) using RL (top), distill the privileged FGP into a depth FGP to predict the teacher’s actions and object position (middle), and deploy DextrAH-G with a state machine in the real world for bin packing (bottom). See Table 3 for details.
@@ 3.1 Geometric Fabrics and Fabric-Guided Policies (FGPs)
where $\mathbf{M}_{f}\in\mathbb{R}^{n\times n}$ is the positive-definite system metric (mass), which captures system prioritization, $\mathbf{f}_{f}\in\mathbb{R}^{n}$ is a nominal path generating geometric force, and $\mathbf{f}_{\pi}(\mathbf{a})\in\mathbb{R}^{n}$ is an additional driving force of some action $\mathbf{a}\in\mathbb{R}^{m}$. $\mathbf{q}_{f},{\dot{\mathbf{q}}}_{f},{\ddot{\mathbf{q}}}_{f}\in\mathbb{R}^{n}$ are the position, velocity, and acceleration of the fabric. This fundamental equation produces an acceleration ${\ddot{\mathbf{q}}}_{f}$, which evolves the fabric state $\mathbf{q}_{f}$ and ${\dot{\mathbf{q}}}_{f}$ over time through numerical integration. One can see that $\mathbf{f}_{\pi}$ influences ${\ddot{\mathbf{q}}}_{f}$, and thereby, the fabric state. For in-depth discussion of geometric fabrics, we refer the reader to prior works [19, 16, 20, 13]. See Appendix B.1 for details on how a geometric fabric connects with a robot.
@@
We use a geometric fabric controller for four primary reasons: 1) avoiding undesirable collisions, 2) creating an inductive bias via the exposed action space that simultaneously guides policy exploration and favorably shapes the overall robot motion, 3) respecting joint constraints (see Appendix B.3), and 4) maintaining the robot posture to promote kinematic manipulability (see Appendix B.4).
@@
Collision Avoidance: Environmental and self-collision avoidance are handled through a geometric fabric term and a forcing fabric term. The geometric term is tuned to dominate the collision avoidance behavior with speed-invariant paths, while the forcing term prevents penetration near the collision boundary. First, we model the geometry of the robot as a collection of spheres and use the forward kinematics of the arm to map the robot configuration to the origin of every sphere attached to the robot body, $\mathbf{x}=\phi_{fk}(\mathbf{q})\in\mathbb{R}^{3}$. We define $\hat{\mathbf{n}}_{i}=\frac{\mathbf{r}_{i}-\mathbf{x}}{\|\mathbf{r}_{i}-\mathbf{x}\|}\in\mathbb{R}^{3}$ as the direction from the sphere point to the closest point on collision body $i$, $\mathbf{r}_{i}\in\mathbb{R}^{3}$ (collision objects or other body spheres) and $\underline{d}_{i}=\max(d_{min},d_{i})\in\mathbb{R}^{+}$ as a lower-bounded distance, where $d_{min}\in\mathbb{R}^{+}$ and $d_{i}\in\mathbb{R}$ is the signed distance between the body sphere and collision body $i$ (Figure 5 for visualization).
The geometric acceleration is ${\ddot{\mathbf{x}}}=k_{g}\|{\dot{\mathbf{x}}}\|^{2}\hat{{\ddot{\mathbf{x}}}}_{b}$ and the forcing acceleration is ${\ddot{\mathbf{x}}}=k_{f}\hat{{\ddot{\mathbf{x}}}}_{b}-b{\dot{\mathbf{x}}}$, where $k_{g},k_{f}\in\mathbb{R}^{+}$ are gains and $b\in\mathbb{R}^{+}$ is a damping scalar. ${\ddot{\mathbf{x}}}_{b}=-\sum_{i}\frac{1}{\underline{d}_{i}}\hat{\mathbf{n}}_{i}$ is a base acceleration response per sphere away from collision, and $\hat{{\ddot{\mathbf{x}}}}_{b}=\frac{{\ddot{\mathbf{x}}}_{b}}{\|{\ddot{\mathbf{x}}}_{b}\|}$ is the normalized base acceleration. The geometric and forcing terms both have metric design $\mathbf{M}=\frac{\beta}{\widetilde{d}^{2}}\hat{\mathbf{M}}_{b}$, where $\beta\in\mathbb{R}^{+}$ is a gain, and $\widetilde{d}=\min_{i}\{\underline{d}_{i}\}$. $\mathbf{M}_{b}=\sum_{i}\frac{s_{i}}{\underline{d}_{i}}\hat{\mathbf{n}}_{i}\otimes\hat{\mathbf{n}}_{i}$ is a base metric response per sphere, and $\hat{\mathbf{M}}_{b}=\frac{\mathbf{M}_{b}}{\|\mathbf{M}_{b}\|}$ is the normalized metric, which maintains its Eigenspectrum (directions of importance). $s_{i}=\frac{1}{2}\tanh(-\alpha_{1}(v_{i}-\alpha_{2})+1)$ is a smooth velocity gate that goes high when this sphere is moving towards collision body $i$, where $\alpha_{1},\alpha_{2}\in\mathbb{R}^{+}$ are gains and $v_{i}=-{\dot{\mathbf{x}}}\cdot\hat{\mathbf{n}}_{i}$ is the signed impact speed that is negative when moving towards collision body $i$ (see Appendix B.2 for details).
@@
Action Space: During grasping, human finger motions exhibit highly correlated movements that do not fully exercise all the degrees of actuation afforded by our hands. These grasps can most broadly be categorized into power and precision grasps [21]. Absent intricate in-hand manipulation, grasping motion can be concisely captured by movement on a lower dimensional manifold. Constraining a policy’s action space to an eigengrasp manifold has been shown to improve grasp policy learning [11]. We create such a manifold for the Allegro hand by retargeting human grasping motion data to the Allegro (see Appendix D for retargeting details) and apply principal component analysis to the motion dataset. Let $\mathbf{A}\in\mathbb{R}^{5\times 16}$ be the first five principal components from PCA and define $\widetilde{\mathbf{A}}=[\mathbf{0},\mathbf{A}]\in\mathbb{R}^{5\times 23}$, then the taskmap from the full robot configuration space to PCA space is $\mathbf{x}=\widetilde{\mathbf{A}}\mathbf{q}\in\mathbb{R}^{5}$. Within this taskmap, we define an attraction fabric term. The metric $\mathbf{M}(\mathbf{x})=m\mathbf{I}$ is a constant isotropic mass, where $m\in\mathbb{R}^{+}$. ${\ddot{\mathbf{x}}}=-k_{a}\tanh(\alpha_{a}\|\mathbf{x}-\mathbf{x}_{pca,target}\|)\frac{\mathbf{x}-\mathbf{x}_{pca,target}}{\|\mathbf{x}-\mathbf{x}_{pca,target}\|}-b{\dot{\mathbf{x}}}$, where $k_{a},\alpha_{a}\in\mathbb{R}^{+}$ are gains and $\mathbf{x}_{pca,target}\in\mathbb{R}^{5}$ is a target position in this space. We use $\mathbf{x}_{pca,target}$ as a 5 dimensional action space for the hand (Figure 7 for example motions).
@@
To coordinate finger control with arm control, we create another action space that controls the pose of the palm. We accomplish this by creating a new taskmap that uses forward kinematics to map to 7 three-dimensional points attached to the palm, stacked into a 21 dimensional space. The attraction fabric in this space is the same as designed for the hand, but scales to the full 21 dimensions with the goal action position $\mathbf{x}_{g}\in\mathbb{R}^{21}$. We create a 6 dimensional action space for the arm that consists of target palm position $\mathbf{x}_{f,target}\in\mathbb{R}^{3}$ and target palm orientation in Euler angles $\mathbf{r}_{f,target}\in\mathbb{R}^{3}$. These quantities are transformed to 3D point targets for all 7 palm-fixed points, i.e., $\mathbf{x}_{g}$, and issued to the fabric. Across the full robot, the action space is 11 dimensional. This action space resulted in faster training and more natural grasping behavior (see Appendix C).
@@ 3.2 Teacher Privileged FGP Training (Reinforcement Learning)
Next, we cast dexterous grasping as a reinforcement learning problem and train a privileged-state teacher policy in simulation to adeptly grasp 140 different objects (additional details in Appendix E). Since the geometric fabric action space ensures that the robot will execute safe and natural behaviors, our reward design centers entirely around fingertip-object contact and lifting the object to a goal. This simplifies reward engineering and reduces reward hacking behavior from the policy.
@@
Environment Modifications for Robust Grasping: RL policies often converge to unnatural solutions that work well in simulation, but fail to transfer to the real world (see Appendix E.13 for frail grasping behavior.) We learn robust grasping behavior by introducing the following environment modifications. Random Wrench Perturbations: We apply random wrenches that move and rotate the object in unpredictable ways (see Appendix E.3). This forces the policy to learn grasps that are robust to exogenous perturbations. Pose Noise: We add uncorrelated and correlated noise to the object pose observation (see Appendix E.4). This gives incentive to learn grasping behavior that opens the hand wider than typically needed when approaching the object to reduce unexpected contact and account for uncertainty in position and geometry. Friction Reduction: We reduce the default coefficient of friction of the object to $\mu=0.7$, mitigating grasping behavior that is overly reliant on friction. Domain Randomization: We employ domain randomization across simulation parameters to learn policies that are robust across a spectrum of dynamics (see Appendix E.5).